# Tiny Time Mixers (TTMs): Fast Pretrained Models for Enhanced Zero/Few-Shot Forecasting of Multivariate Time Series

### 11 Jan 2024 v2

@@ 3.2 Fine-tuning Workflow
Zero-shot Few-shot (5%) $fl$ TTM TTM GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 96 0.365 0.366 0.543 0.557 0.554 0.892 0.547 0.593 0.681 192 0.393 0.391 0.748 0.711 0.673 0.940 0.720 0.652 0.725 336 0.415 0.421 0.754 0.816 0.678 0.945 0.984 0.731 0.761 720 0.538 - - - - - - - - ETTH2 96 0.285 0.282 0.376 0.401 0.348 0.409 0.442 0.390 0.428 192 0.341 0.338 0.418 0.452 0.419 0.483 0.617 0.457 0.496 336 0.383 0.383 0.408 0.464 0.389 0.499 1.424 0.477 0.486 720 0.441 - - - - - - - - ETTM1 96 0.413 0.359 0.386 0.399 0.361 0.606 0.332 0.628 0.726 192 0.476 0.402 0.440 0.441 0.411 0.681 0.358 0.666 0.750 336 0.553 0.424 0.485 0.499 0.467 0.786 0.402 0.807 0.851 720 0.737 0.575 0.577 0.767 0.677 0.796 0.511 0.822 0.857 ETTM2 96 0.187 0.174 0.199 0.206 0.200 0.220 0.236 0.229 0.232 192 0.261 0.240 0.256 0.264 0.265 0.311 0.306 0.394 0.291 336 0.323 0.299 0.318 0.334 0.314 0.338 0.380 0.378 0.478 720 0.436 0.407 0.46 0.454 0.410 0.509 0.674 0.523 0.553 Weather 96 0.154 0.152 0.175 0.171 0.188 0.207 0.184 0.229 0.227 192 0.203 0.198 0.227 0.230 0.234 0.272 0.228 0.265 0.278 336 0.256 0.250 0.286 0.294 0.287 0.313 0.279 0.353 0.351 720 0.329 0.326 0.366 0.384 0.365 0.400 0.364 0.391 0.387 Electricity 96 0.169 0.142 0.143 0.145 0.147 0.315 0.150 0.235 0.297 192 0.196 0.162 0.159 0.163 0.172 0.318 0.163 0.247 0.308 336 0.209 0.184 0.179 0.175 0.190 0.340 0.175 0.267 0.354 720 0.264 0.228 0.233 0.219 0.280 0.635 0.219 0.318 0.426 Traffic 96 0.518 0.401 0.419 0.404 0.408 0.854 0.427 0.670 0.795 192 0.548 0.425 0.434 0.412 0.421 0.894 0.447 0.653 0.837 336 0.55 0.437 0.449 0.439 0.477 0.853 0.478 0.707 0.867 720 0.605 - - - - - - - - TTM Zero-shot % IMP 1% 3% 2% 31% 6% 25% 32% TTM 5% Few-shot % IMP 12% 14% 12% 38% 17% 32% 38% Table 1: Zero-shot and Few-shot 5% performance of TTM and all SOTA models on D1 datasets for varying forecast lengths ($fl$). Bold and underscore denote the best and the second best. For longer $fl$ (720), some datasets fail to generate fine-tuning samples with just 5% of the data, denoted as ‘-’. The last two rows show relative improvement (IMP) of TTM zero-shot and few-shot w.r.t. SOTA few-shot results.
@@
Metric Data GPT4TS TTM (Ours) IMP (X) FT NPARAMS (M) Weather 3.2 0.29 (14X) Electricity 3.2 0.29 Traffic 5.5 0.29 TL NPARAMS (M) Weather 84.3 0.8 (106X) Electricity 84.3 0.8 Traffic 86.6 0.8 EPOCH TIME (s) Weather 37 1 (65X) Electricity 241 4 Traffic 782 8 MAX MEMORY (GB) Weather 21.9 1.06 (27X) Electricity 39.7 2.02 Traffic 53.7 1.34 TEST TIME (s) Weather 52.53 1.66 (54X) Electricity 353.4 8.07 Traffic 1163.9 13.5 Table 2: Computational Improvement in few-shot 10% for $fl$=96 using 1 A100 GPU. nX denotes n times average improvement across datasets (Avg. IMP). 10% 25% 50% 75% 100% IMP TTM 0.422 0.421 0.413 0.402 0.398 - SimMTM 0.591 0.535 0.491 0.466 0.415 17% Ti-MAE 0.660 0.594 0.55 0.475 0.466 24% TST 0.783 0.641 0.525 0.516 0.469 28% LaST 0.645 0.610 0.540 0.479 0.443 23% TF-C 0.799 0.736 0.731 0.697 0.635 43% CoST 0.784 0.624 0.540 0.494 0.428 26% TS2Vec 0.655 0.632 0.599 0.577 0.517 31%
@@
Figure 2: Cross transfer learning improvement (IMP) for self-supervised pretraining methods in various few-shot settings (10%,25%,50%,75%,100%).
@@ 4.1 Experimental Setting
Datasets: Pretraining employs a subset of the Monash data hub Godahewa et al. (2021) of size $\sim 240$M samples. We specifically exclude datasets that consist of yearly, monthly, and daily series as they do not possess sufficient length for the long-term forecasting task. Moreover, we remove all the datasets that we utilize for evaluation. This includes weather, electricity, and traffic datasets. For zero/few/full-shot evaluation we consider seven datasets (D1): ETTH1, ETTH2, ETTM1, ETTM2, Weather, Electricity (ECL), Traffic as popularly used in most prior SOTA works Zhou et al. (2021); Nie et al. (2022). Since these datasets do not contain any exogenous variables nor exhibit cross-channel correlation benefits, we incorporate four other datasets (D2) for separately validating the efficacy of the decoder channel mixing and exogenous mixer module: bike sharing (BS) Fanaee-T (2013), carbon capture plant (CC) Jablonka et al. (2023), and two more datasets from Business-IT observability domain Biz (2023): Application (APP) and Service (SER). Details are provided in the Appendix.
@@
SOTA Bechmarks: We benchmark TTM model with the latest publicly available SOTA models for multivariate forecasting categorized as follows: [label=()] 1. LLM-based TS pretrained models:GPT4TS Zhou et al. (2023), LLMTime Gruver et al. (2023) 2. Self-supervised pretrained models: SimMTM Dong et al. (2023),Ti-MAE Li et al. (2023), TST Zerveas et al. (2021), LaST Wang et al. (2022), TF-C Zhang et al. (2022), CoST Woo et al. (2022) and Ts2Vec Yue et al. (2022) 3. TS transformer models:PatchTST Nie et al. (2022), FEDFormer Zhou et al. (2022), Autoformer Wu et al. (2021) 4. Other SOTA models:TSMixer Ekambaram et al. (2023), DLinear Zeng et al. (2022) and TimesNet Wu et al. (2022)
@@
Zero-shot Few-shot (10%) Data TTM (Ours) TTM (Ours) GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 0.428 0.422 0.590 0.633 0.578 0.869 0.691 0.639 0.702 ETTH2 0.362 0.360 0.397 0.415 0.370 0.479 0.605 0.466 0.488 ETTM1 0.545 0.458 0.464 0.501 0.441 0.677 0.411 0.722 0.802 ETTM2 0.302 0.278 0.293 0.296 0.282 0.320 0.316 0.463 1.342 Electricity 0.210 0.180 0.176 0.176 0.180 0.323 0.180 0.346 0.431 Traffic 0.555 0.428 0.440 0.430 0.429 0.951 0.446 0.663 0.749 Weather 0.236 0.227 0.238 0.242 0.236 0.279 0.241 0.284 0.300 TTM Zero-shot improvements -4% -2% -7% 27% 2% 27% 39% TTM Few-shot (10%) improvements 7% 9% 4% 34% 13% 34% 45% Table 3: Zero-shot and Few-shot 10% performance of TTM and all SOTA models on seven datasets averaged across 4 forecast lengths. The numbers with bold and underscore denote the best the second best results respectively. The full table is in Appendix.
@@
Data $fl$ Zero-shot LLMTime TTM Traffic 96 0.340 0.195 192 0.526 0.367 Weather 96 0.107 0.031 192 0.062 0.067 Electricity 96 0.609 0.238 192 0.960 0.488 ETTM2 96 0.167 0.164 192 0.198 0.228 TTM IMP (%) 29% Table 4: Comparison between TTM and LLMTime in zero-shot setting.
@@
TTM Model Details: By default, the following data and model configuration is used for pretraining: input sequence length $sl$ = 512, patch length $pl$ = 64, stride $s$ = 64, number of patches $n=512/64=8$, forecast length $fl\in\{96,192,336,720\}$, number of levels $L$ = 6, number of TTM blocks per level $M$ = 2, batch size $b$ = 3000, number of epochs $ep$ = 25, and dropout $do$ = 0.2. TSMixer specific hyperparameters include feature scaler $fs$ = 3, hidden feature size $hf$ = $fs*pl=192$, expansion feature size $ef=hf*2=384$. Please note that, $hf$ and $n$ will change across TTM blocks based on the adaptive patching strategy, and the decoder works in channel-independent mode during pretraining. Resolution prefix tuning is disabled by default and enabled only for shorter context length (as explained in Table. 9). Training is performed in a distributed fashion with 50 CPUs and 6 NVIDIA A100 GPUs with 40 GB GPU memory. We build a pretrained model using Monash data for each specified $fl$. Since pretrained models are very small, they can be trained quickly in a few hours (4-8 hrs based on $fl$ length), as opposed to several days/weeks in standard approaches. For few/full-shot fine-tuning, all the above pretraining parameters remain the same except the following. Based on the target data characteristics, the decoder channel-mixing and/or exogenous mixer block is enabled/disabled (both enabled for D2 datasets). In addition, head dropout changes to 0.7 for smaller ETT datasets and 0.2 otherwise and batch size ranges between 8-64. All the hyperparameters are adjusted based on the validation performance, and the final test results are reported in the paper.
Data BS CC APP SER % IMP
@@
Table 5: Effect of Decoder Mixing and Exog Fusion. Results are reported using $(sl,fl)$ with values of $(512,96)$ for BS and $(96,24)$ for other datasets.
@@ 4.2 TTM’s Zero/Few-shot Performance
Table 1 compares the performance of TTM model in zero-shot and few-shot (5%) settings. Baseline results are reported from  Zhou et al. (2023) as we use the same few-shot data filtering strategy as followed in that paper. Note that for zero-shot performance, the TTM model pretrained on the Monash data is directly evaluated on the test set for every dataset. We report the test MSE for multiple forecast horizons varying from 96 to 720. The 5% few-shot TTM outperforms the SOTAs in most of the cases with significant accuracy gains (12-38%). An even more impressive observation is that the TTM in a zero-shot setting is also able to outperform most of the SOTAs which are trained on 5% of the target data. This observation establishes the generalization ability of the pretrained TTM model on the target datasets. Likewise, Table 3 shows the 10% few-shot performance of the TTM, where we outperform all the existing SOTAs by 4-45% accuracy gains. In addition, TTM zero-shot also beats many SOTAs (not all) with 10% training highlighting the effectiveness of our approach.
Additionally, we conduct a comparison between TTM and the LLMTime model which is explicitly designed for zero-shot setting. Since these models are based on LLaMA and are massive in size, the authors used only the last window of the standard test-set for faster evaluation, as opposed to using all the windows in the full test-set . Hence, we compare them separately in Table 4 based on the same datasets, $fl$, and test-set as reported in  Gruver et al. (2023). In this evaluation, we outperform LLMTime by a substantial margin of 29%. In addition, there are alternative pretraining approaches like masked modeling and contrastive learning techniques that may not offer universal transfer learning capabilities across all datasets (like TTM), but they excel in enabling cross-transfer learning between two datasets when carefully selected. Table 2 illustrates the cross-transfer learning from ETTH2 to ETTH1 for these models in various few-shot settings (as reported in  Dong et al. (2023)). Notably, TTM, with no specific cross-data selection, outperforms all popular SOTAs, including the latest SimMTM, by 17-43%. Thus, TTM significantly beats all the existing SOTAs including the recent popular transfer learning benchmarks based on LLMs. The notable part here is that we get these substantial accuracy gains with significantly reduced compute/runtime requirements as explained next.
@@ 5 Epochs 50 Epochs 5 Epochs 50 Epochs
Table 6: Effect of pretraining (PT) over Randomly Initialized (RI) weights, averaged across 4 values of $fl$. Full table in Appendix.
Data Zero-shot Few-shot (10%) Vanilla Adaptive Patching Vanilla Adaptive Patching ETTH1 0.501 0.428 0.478 0.422 ETTH2 0.375 0.362 0.372 0.360 ETTM1 0.577 0.545 0.463 0.458 ETTM2 0.310 0.302 0.272 0.278 Electricity 0.215 0.210 0.178 0.180 Traffic 0.556 0.555 0.431 0.428 Weather 0.239 0.236 0.230 0.227 IMP (%) 4% 2% Table 7: Effect of adaptive Patching, averaged across 4 values of $fl$.
Data Zero-shot Original Data With Downsample ETTH1 0.603 0.379 ETTH2 0.351 0.313 ETTM1 0.852 0.444 ETTM2 0.261 0.224 Electricity 0.424 0.182 Traffic 0.898 0.533 Weather 0.186 0.178 IMP (%) 30% Table 8: Effect of downsampling, averaged across $fl$ 96 and 192.
Data $HL$ = 96 $HL$ = 512 NP P NP P ETTH1 0.373 0.358 0.365 0.360 ETTH2 0.180 0.179 0.285 0.280 ETTM1 0.559 0.387 0.413 0.384 ETTM2 0.127 0.108 0.187 0.194 Weather 0.103 0.103 0.154 0.160 Electricity 0.208 0.201 0.169 0.175 Traffic 0.754 0.740 0.518 0.518 IMP (%) 8% 0% Table 9: Resolution Prefix Tuning: With (P) vs Without(NP)

### 17 Jan 2024 v3

@@ 3.2 Fine-tuning Workflow
Zero-shot Few-shot (5%) $fl$ TTM TTM GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 96 0.365 0.366 0.543 0.557 0.554 0.892 0.547 0.593 0.681 192 0.393 0.391 0.748 0.711 0.673 0.940 0.720 0.652 0.725 336 0.415 0.421 0.754 0.816 0.678 0.945 0.984 0.731 0.761 720 0.538 - - - - - - - - ETTH2 96 0.285 0.282 0.376 0.401 0.348 0.409 0.442 0.390 0.428 192 0.341 0.338 0.418 0.452 0.419 0.483 0.617 0.457 0.496 336 0.383 0.383 0.408 0.464 0.389 0.499 1.424 0.477 0.486 720 0.441 - - - - - - - - ETTM1 96 0.413 0.359 0.386 0.399 0.361 0.606 0.332 0.628 0.726 192 0.476 0.402 0.440 0.441 0.411 0.681 0.358 0.666 0.750 336 0.553 0.424 0.485 0.499 0.467 0.786 0.402 0.807 0.851 720 0.737 0.575 0.577 0.767 0.677 0.796 0.511 0.822 0.857 ETTM2 96 0.187 0.174 0.199 0.206 0.200 0.220 0.236 0.229 0.232 192 0.261 0.240 0.256 0.264 0.265 0.311 0.306 0.394 0.291 336 0.323 0.299 0.318 0.334 0.314 0.338 0.380 0.378 0.478 720 0.436 0.407 0.46 0.454 0.410 0.509 0.674 0.523 0.553 Weather 96 0.154 0.152 0.175 0.171 0.188 0.207 0.184 0.229 0.227 192 0.203 0.198 0.227 0.230 0.234 0.272 0.228 0.265 0.278 336 0.256 0.250 0.286 0.294 0.287 0.313 0.279 0.353 0.351 720 0.329 0.326 0.366 0.384 0.365 0.400 0.364 0.391 0.387 Electricity 96 0.169 0.142 0.143 0.145 0.147 0.315 0.150 0.235 0.297 192 0.196 0.162 0.159 0.163 0.172 0.318 0.163 0.247 0.308 336 0.209 0.184 0.179 0.175 0.190 0.340 0.175 0.267 0.354 720 0.264 0.228 0.233 0.219 0.280 0.635 0.219 0.318 0.426 Traffic 96 0.518 0.401 0.419 0.404 0.408 0.854 0.427 0.670 0.795 192 0.548 0.425 0.434 0.412 0.421 0.894 0.447 0.653 0.837 336 0.55 0.437 0.449 0.439 0.477 0.853 0.478 0.707 0.867 720 0.605 - - - - - - - - TTM Zero-shot % IMP 1% 3% 2% 31% 6% 25% 32% TTM 5% Few-shot % IMP 12% 14% 12% 38% 17% 32% 38% Table 1: Zero-shot and Few-shot 5% performance (MSE) of TTM and all SOTA models on D1 datasets for varying forecast lengths ($fl$). Bold and underscore denote the best and the second best. For longer $fl$ (720), some datasets fail to generate fine-tuning samples with just 5% of the data, denoted as ‘-’. The last two rows show relative improvement (IMP) of TTM zero-shot and few-shot w.r.t. SOTA few-shot results.
@@
Metric Data GPT4TS TTM (Ours) IMP (X) FT NPARAMS (M) Weather 3.2 0.29 (14X) Electricity 3.2 0.29 Traffic 5.5 0.29 TL NPARAMS (M) Weather 84.3 0.8 (106X) Electricity 84.3 0.8 Traffic 86.6 0.8 EPOCH TIME (s) Weather 37 1 (65X) Electricity 241 4 Traffic 782 8 MAX MEMORY (GB) Weather 21.9 1.06 (27X) Electricity 39.7 2.02 Traffic 53.7 1.34 TEST TIME (s) Weather 52.53 1.66 (54X) Electricity 353.4 8.07 Traffic 1163.9 13.5 Table 2: Computational Improvement in few-shot 10% for $fl$=96 using 1 A100 GPU. nX denotes n times average improvement across datasets (IMP). 10% 25% 50% 75% 100% IMP TTM 0.422 0.421 0.413 0.402 0.398 - SimMTM 0.591 0.535 0.491 0.466 0.415 17% Ti-MAE 0.660 0.594 0.55 0.475 0.466 24% TST 0.783 0.641 0.525 0.516 0.469 28% LaST 0.645 0.610 0.540 0.479 0.443 23% TF-C 0.799 0.736 0.731 0.697 0.635 43% CoST 0.784 0.624 0.540 0.494 0.428 26% TS2Vec 0.655 0.632 0.599 0.577 0.517 31%
@@
Figure 2: Cross transfer learning MSE improvement (IMP) for self-supervised pre-training methods in various few-shot settings (10%,25%,50%,75%,100%).
@@ 4.1 Experimental Setting
Datasets: Pre-training employs a subset of the Monash data hub Godahewa et al. (2021) of size $\sim 244$M samples. We specifically exclude datasets (like yearly, monthly) as they do not possess sufficient length for the long-term forecasting task. Moreover, we remove all the datasets that we utilize for evaluation (i.e., weather, electricity, and traffic). For zero/few-shot evaluation we consider seven public datasets (D1): ETTH1, ETTH2, ETTM1, ETTM2, Weather, Electricity & Traffic as popularly used in most prior SOTA works Zhou et al. (2021); Nie et al. (2023). Since these datasets do not contain any exogenous variables nor exhibit cross-channel correlation benefits, we incorporate four other datasets (D2) for separately validating the efficacy of the decoder channel mixing and exogenous mixer module: bike sharing (BS) Fanaee-T (2013), carbon capture plant (CC) Jablonka et al. (2023), and 2 more datasets from Biz-IT observability domain ITB (2023): Application (APP) and Service (SER). Refer Appendix for full data details.
@@
SOTA Bechmarks: We benchmark TTM with the latest public SOTA forecasting models categorized as follows: [label=()] 1. LLM-based TS pre-trained models:GPT4TS Zhou et al. (2023), LLMTime Gruver et al. (2023) 2. Self-supervised pre-trained models: SimMTM Dong et al. (2023),Ti-MAE Li et al. (2023), TST Zerveas et al. (2021), LaST Wang et al. (2022), TF-C Zhang et al. (2022), CoST Woo et al. (2022) and Ts2Vec Yue et al. (2022) 3. TS transformer models:PatchTST Nie et al. (2023), FEDFormer Zhou et al. (2022), Autoformer Wu et al. (2021) 4. Other SOTA models:TSMixer Ekambaram et al. (2023), DLinear Zeng et al. (2022) and TimesNet Wu et al. (2022)
@@
Zero-shot Few-shot (10%) Data TTM (Ours) TTM (Ours) GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 0.428 0.422 0.590 0.633 0.578 0.869 0.691 0.639 0.702 ETTH2 0.362 0.360 0.397 0.415 0.370 0.479 0.605 0.466 0.488 ETTM1 0.545 0.458 0.464 0.501 0.441 0.677 0.411 0.722 0.802 ETTM2 0.302 0.278 0.293 0.296 0.282 0.320 0.316 0.463 1.342 Electricity 0.210 0.180 0.176 0.176 0.180 0.323 0.180 0.346 0.431 Traffic 0.555 0.428 0.440 0.430 0.429 0.951 0.446 0.663 0.749 Weather 0.236 0.227 0.238 0.242 0.236 0.279 0.241 0.284 0.300 TTM Zero-shot improvements -4% -2% -7% 27% 2% 27% 39% TTM Few-shot (10%) improvements 7% 9% 4% 34% 13% 34% 45% Table 3: Zero-shot and Few-shot 10% performance (MSE) of TTM over all SOTA models on D1 datasets, averaged across 4 forecast lengths. The numbers with bold and underscore denote the best and the second best results respectively. The full table is in Appendix.
@@
Data $fl$ Zero-shot LLMTime TTM Traffic 96 0.340 0.195 192 0.526 0.367 Weather 96 0.107 0.031 192 0.062 0.067 Electricity 96 0.609 0.238 192 0.960 0.488 ETTM2 96 0.167 0.164 192 0.198 0.228 TTM IMP (%) 29% Table 4: TTM vs LLMTime MSE Improvement (IMP) in zero-shot setting.
@@
TTM Model Details: For our experiments, we build 5 pre-trained models for the following $sl$,$fl$ configuration: (512,96), (512,192), (512, 336), (512,720) and (96,24). Based on the $sl$ and $fl$ requirement of the target dataset, a suitable pre-trained model is selected to initialize the weights accordingly for fine-tuning. Since our pre-trained models are very small, they can be trained quickly in a few hours (4-8 hrs), as opposed to several days/weeks in standard approaches. Hence, pre-training multiple TTM models is no longer a practical constraint. Pre-training is performed in a distributed fashion with 50 CPUs and 6 A100 GPUs with 40 GB GPU memory while fine-tuning uses only 1 GPU. Other model configurations are as follows: patch length $pl$ = 64 (for $sl$ 512) and 8 (for $sl$ 96), stride $s$ = $pl$, number of levels $L$ = 6, number of TTM blocks per level $M$ = 2, number of decoder layers = 2, batch size $b$ = 3K, number of epochs $ep$ = 20, and dropout $do$ = 0.2. TSMixer specific hyperparameters include feature scaler $fs$ = 3, hidden feature size $hf$ = $fs*pl$, expansion feature size $ef=hf*2$. Resolution prefix tuning is disabled by default and enabled only for shorter context lengths (as explained in Table. 9). Decoder channel-mixing and exogenous mixer blocks are disabled during pre-training and enabled for D2 datasets during fine-tuning. All other model parameters remain the same across pre-training and fine-tuning except for a few parameters like dropout and batch size that can be adjusted based on the target data. Hyperparameters are selected based on the validation performance, and the test results are reported. For full implementation details, refer to the Appendix. MSE is used as the default evaluation metric.
BS CC APP SER % IMP
@@
Table 5: Effect of Decoder Mixing and Exog Fusion. MSE Results are reported using $(sl,fl)$ with values of $(512,96)$ for BS dataset and $(96,24)$ for other D2 datasets.
@@ 4.2 TTM’s Zero/Few-shot Performance
Table 1 compares the performance of TTM model in zero-shot and few-shot (5%) settings across multiple $fl$s. Baseline results are reported from  Zhou et al. (2023) as we use the same few-shot data filtering strategy as followed in that paper. Note that for zero-shot performance, the pre-trained TTM model is directly evaluated on the test set. The 5% few-shot TTM outperforms the SOTAs in most of the cases with significant accuracy gains (12-38%). An even more impressive observation is that the TTM in zero-shot setting is also able to outperform most of the SOTAs which are trained on 5% of the target data. This observation establishes the generalization ability of the pre-trained TTM model on the target datasets. Likewise, Table 3 shows the 10% few-shot performance of the TTM, where we outperform all the existing SOTAs by 4-45% accuracy gains. In addition, TTM zero-shot also beats many SOTAs (not all) with 10% training highlighting the effectiveness of our approach.
Additionally, we conduct a comparison between TTM and the LLMTime model which is explicitly designed for zero-shot setting. Since these models are based on LLaMA and are massive in size, the authors used only the last window of the standard test-set for faster evaluation, as opposed to using all the windows in the full test-set . Hence, we compare them separately in Table 4 based on the same datasets, $fl$, and test-set as reported in  Gruver et al. (2023). In this evaluation, we outperform LLMTime by a substantial margin of 29%. In addition, there are alternative pre-training approaches like masked modeling and contrastive learning techniques that may not offer universal transfer learning capabilities across all datasets (like TTM), but they excel in enabling cross-transfer learning between two datasets when carefully selected. Table 2 illustrates the cross-transfer learning from ETTH2 to ETTH1 for these models in various few-shot settings (as reported in  Dong et al. (2023)). Notably, TTM, with no specific cross-data selection, outperforms all popular SOTAs, including the latest SimMTM, by 17-43%. Thus, TTM significantly beats all the existing SOTAs including the recent popular transfer learning benchmarks based on LLMs. Notably, we achieve these accuracy gains with significantly reduced compute requirements, as elaborated next.
@@ 5 Epochs 50 Epochs 5 Epochs 50 Epochs
Table 6: Improvement (IMP) of Pre-Training (PT) over Randomly Initialized (RI) weights. MSE averaged across all $fl$s.
Data Zero-shot Few-shot (10%) Vanilla Adaptive Patching Vanilla Adaptive Patching ETTH1 0.501 0.428 0.478 0.422 ETTH2 0.375 0.362 0.372 0.360 ETTM1 0.577 0.545 0.463 0.458 ETTM2 0.310 0.302 0.272 0.278 Electricity 0.215 0.210 0.178 0.180 Traffic 0.556 0.555 0.431 0.428 Weather 0.239 0.236 0.230 0.227 IMP (%) 4% 2% Table 7: Effect of adaptive Patching, MSE averaged across 4 values of $fl$
Data Zero-shot Original Data With Downsample ETTH1 0.603 0.379 ETTH2 0.351 0.313 ETTM1 0.852 0.444 ETTM2 0.261 0.224 Electricity 0.424 0.182 Traffic 0.898 0.533 Weather 0.186 0.178 IMP (%) 30% Table 8: Downsampling effect, MSE averaged across $fl$s: 96, 192
Data $HL$ = 96 $HL$ = 512 NP P NP P ETTH1 0.373 0.358 0.365 0.360 ETTH2 0.180 0.179 0.285 0.280 ETTM1 0.559 0.387 0.413 0.384 ETTM2 0.127 0.108 0.187 0.194 Weather 0.103 0.103 0.154 0.160 Electricity 0.208 0.201 0.169 0.175 Traffic 0.754 0.740 0.518 0.518 IMP (%) 8% 0% Table 9: Resolution Prefix Tuning:With (P) vs Without(NP). MSE reported